OpenStudy (anonymous):

Simplify 5 + 2{x - 4[3x + 7(2 - x)]}

OpenStudy (callisto):

5 + 2{x - 4[3x + 7(2 - x)]} = 5 + 2{x - 4[3x + 7(2) - 7x]} = 5 + 2{x - 4[3x - 7x + 14]} = 5 + 2{x - 4[-4x + 14]} = 5 + 2{x - 4(-4x) + (-4)14]} = 5 + 2{x +16x - 56} = 5 + 2{17x - 56} = 5 + 2(17x) - 2(56) = ... ?

OpenStudy (anonymous):

i need an asnwer to to question not bs

OpenStudy (callisto):

Simplify it and get your answer

OpenStudy (anonymous):

34x + 107?

OpenStudy (callisto):

34x => correct. +107 => incorrect. What is - 2(56)?

OpenStudy (anonymous):

is it 34x - 107

OpenStudy (anonymous):

it is neg itive right

OpenStudy (callisto):

Yes!

OpenStudy (anonymous):

thanks

OpenStudy (callisto):

Welcome